OpenStudy (anonymous):

lim (4x-x^2)/(2-sqrtx) x->4 How is it 16?

OpenStudy (anonymous):

factor and cancel, or multiply by the conjugate your choice

OpenStudy (anonymous):

Alright satellite take it away :P .

OpenStudy (luigi0210):

Try multiply by the conjugate

OpenStudy (anonymous):

I rationalized personally.

OpenStudy (anonymous):

i pick factoring

OpenStudy (luigi0210):

Yea.. what Satellite said

OpenStudy (anonymous):

damn typo \[\frac{x(4-x)}{2-\sqrt{x}}=\frac{x(2-\sqrt{x})(2+\sqrt{x})}{2-\sqrt{x}}\]

OpenStudy (anonymous):

Was about to say :P .

OpenStudy (anonymous):

cancel, replace \(x\) by \(16\)

OpenStudy (anonymous):

Replace x with 4 not 16 :P .

OpenStudy (anonymous):

Oh alright thanks a lot! Sometimes I get stuck on little things.

OpenStudy (anonymous):

you can do what @Luigi0210 said too try it and see multiply top and bottom by \(2+\sqrt{x}\) leave the numerator in factored form, cancel the \(4-x\) top and bottom

OpenStudy (anonymous):

oh right, replace \(x\) by \(4\) not \(16\)

OpenStudy (anonymous):

@Dido525 keeping me honest here ...

OpenStudy (anonymous):

I try :) .

OpenStudy (anonymous):

L'hospital's rule would work here too I think.